OpenStudy (anonymous):

Find the first 5 terms of the sequence: an = (-1/2)^n - 1 a) -1, -1/2, -1/4, -1/8, -1/16 b) -1, 1/2, -1/4, 1/8, -1/16 c) 1, -1/2, 1/4, -1/8, 1/16 d) 1, 1/2, 1/4, 1/8, 1/16

OpenStudy (anonymous):

When n = 1, an = 1 When n = 2, an = -1/2 When n = 3, an = 1/4 .......